Suppose that the Treasury bill rate is 6% rather than 3%, as we assumed in Table 12.1 but that the expected return on the market is still 10%. Use the betas in that table to answer the following questions.


a. Calculate the expected return from Pfizer. (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)

b. What is the highest expected return offered by one of these stocks? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)


c. What is the lowest expected return offered by one of these stocks? (Do not round intermediate calculations. Enter your answer as a percent rounded to 1 decimal place.)

TABLE 12.1 Betas for selected common stocks, January 2011-December 2015 Beta U.S. Steel 1.85 Ford General Electric Monsanto Boelng Unlon Paciflc Alphabet ExxonMobll Amazon Intel Pfizer Starbucks IBM McDonalds Coca-Cola Campbell Soup Walmart Newmont Mining PG&E 1.31 1.20 1.19 1.01 1.00 0.96 0.94 0.93 0.91 0.90 0.79 0.59 0.51 0.49 0.47 0.26 0.24 0.23 NoteBetas are calculated from 5 years of monthly data.
Answer #1

risk free rate=6%

return on the market = 10%

Expected returns of all the stocks calculated based on the above formula:

Stock beta risk free rate(%) market return Expected return
US Steel 1.85 6 10 13.4
Disney 1.42 6 10 11.7
Ford 1.31 6 10 11.2
General Electric 1.2 6 10 10.8
Monsanto 1.19 6 10 10.8
Boeing 1.01 6 10 10.0
Union Pacific 1 6 10 10.0
Alphabet 0.96 6 10 9.8
ExxonMobol 0.94 6 10 9.8
Amazon 0.93 6 10 9.7
Intel 0.91 6 10 9.6
Pfizer 0.9 6 10 9.6
Starbucks 0.79 6 10 9.2
IBM 0.59 6 10 8.4
McDonald's 0.51 6 10 8.0
Coca Cola 0.49 6 10 8.0
Campbell Soup 0.47 6 10 7.9
Walmart 0.26 6 10 7.0
Newmont Mining 0.24 6 10 7.0
PG&E 0.23 6 10 6.9

a) expected return from Pfizer= 6%+0.90*(10%-6%)= 9.6%

b) highest expected return is offered by US Steel

expected return of US Steel =6% + 1.85*(10%-6%)= 13.4%

c) Lowest expected return is offered by PG&E

expected return of PG&E =6% + 0.23*(10%-6%)= 6.9%